# REALM: An MLLM-Agent Framework for Open World 3D Reasoning Segmentation and Editing on Gaussian Splatting

### v1

@@ 1 Introduction
  • We propose MLLM-Based Instance Segmenter (LMSeg) that performs image-level reasoning segmentation using MLLM and infers the corresponding Gaussian identity based on the 3D feature field. To produce high-quality 3D object masks, we propose Global-to-Local Spatial Grouding (GLSpaG), which aggregates image-level reasoning segmentations in a global-to-local manner.
@@ 3.4 Global-to-Local Spatial Grounding (Local)
Figure 8: Annotation prompt for REALM3D dataset. We use MLLM to annotate implicit prompts for REALM3D dataset according to the layout of input images.
@@ 4.1 Experimental Settings
(2) REALM3D dataset: To facilitate future research, we introduce REALM3D, a dataset specifically designed to evaluate 3D reasoning segmentation. REALM3D comprises 100+ 3D scenes captured in multiview images, along with 3D point clouds and camera poses generated by VGGT [34]. We annotate 1k+ prompt–mask pairs using Qwen2.5-VL and SAM, covering diverse forms of implicit and explicit prompts (as shown in Fig. 8). REALM3D can be used to evaluate the robustness of models across diverse applications. We provide details of REALM3D in the supplementary materials.
Baselines and metrics. We compare REALM against previous state-of-the-art methods for open-vocabulary 3D segmentation, including GS-Group [35], Gaga [21], and GAGS [24]. We report the mIoU and mBIoU following previous works  [31, 25, 13, 24] to quantitatively exanimate the accuracy of 3D reasoning segmentation results.
@@
Methods mIoU mBIoU Time (s)
Global Reasoning 0.89 0.88 20.21
+Local Reasoning 0.89 0.88 79.68
+Local Refinement 0.95 0.94 83.35
(a) Performance and efficiency of each component in GLSpaG. If faster inference is desired, the framework can operate using only the global grouping stage.
@@
(b) Global camera sampling. K-means view clustering and Top-K ID selection play a crucial role in the effectiveness of the global camera sampling process.
@@
(d) K-means Clusters. Both insufficient and excessive clustering can affect the results of multi-view reasoning.
Value of $N^{\text{global}}$ mIoU mBIoU Speed (s)
$N^{\text{global}}=1$ 0.43 0.42 65.36
$N^{\text{global}}=2$ 0.43 0.42 70.40
$N^{\text{global}}=8$ 0.95 0.94 83.35
$N^{\text{global}}=16$ 0.95 0.94 89.50
(e) Number of global cameras. Too few views harm accuracy; too many views may slow down inference speed.
@@
(f) Local refinement steps. We evaluate the sensitivity of REALM to the local refinement steps. Excessive finetuning may lead to overfitting and degradation.
@@ 4.2 Main Results
Qualitative Comparisons. Previous methods enable 3D localization by leveraging the language understanding capabilities of CLIP [26] or Grounded-SAM [28]. While these approaches offer basic open-vocabulary querying capabilities, they lack the ability to perform reasoning over implicit instructions. We visualize the performance between REALM and baselines under different type of implicit queries. The results are presented in Fig. 5.
@@
Quantitative Comparisons. We quantitatively evaluate the performance of REALM on both implicit and explicit queries. A subset of results is presented in Tab. 1 and LABEL:tab:realm3d. More results can be found in the supplementary materials.
@@
Language-Driven 3D Editing. With accurate 3D object localization, REALM enables precise and fine-grained scene editing without disturbing surrounding objects. As shown in Fig. 7, our model supports a variety of 3D editing tasks, including object removal, replacement, and stylization. REALM modifies the scene without interfering with adjacent content, ensuring faithful preservation of occlusion relationships. Even in tasks involving large-scale appearance changes, such as stylization, REALM effectively isolates the target object while leaving surrounding regions unaffected.
@@ 4.3 Ablation Study
Each component of GLSpaG. As shown in Tab. 2(a), we evaluate both the performance and runtime of the model after completing each stage of GLSpaG. For faster inference, the framework can operate using only the global grouping stage. The local grouping stage, while incurring additional inference time, yields more precise results.
@@
Sensitivity Analysis of Hyper-parameters. The results can be found in the supplementary materials.

### v2

@@ 1 Introduction
  • In REALM, MLLM-Based Instance Segmenter is proposed to perform image-level reasoning segmentation and infer the corresponding Gaussian identity. To produce accurate 3D object masks, we propose Global-to-Local Spatial Grouding, which aggregates image-level reasoning segmentations in a global-to-local manner.
@@ 3.4 Global-to-Local Spatial Grounding (Local)
Num. of Scenes Prompt-Mask Pairs Implicit Prompts
LERF [13] 5 36
3DOVS [20] 10 150
REALM3D 100 1444
Table 2: Statistics of 3D segmentation benchmark. Compared to existing benchmarks, our REALM3D contains more scenes and prompt-mask pairs. Additionally, REALM3D provides implicit prompts for each mask annotation.
Figure 6: Examples in REALM3D benchmark. We use MLLM [3] and SAM [16] to annotate over 1K prompt–mask pairs, enabling quantitative evaluation on implicit queries.
@@ 4.1 Experimental Settings
(2) REALM3D dataset: To facilitate future research, we introduce REALM3D, a dataset specifically designed to evaluate 3D reasoning segmentation. REALM3D comprises 100+ 3D scenes captured in multiview images, along with 3D point clouds and camera poses generated by VGGT [34]. We annotate 1k+ prompt–mask pairs using Qwen2.5-VL and SAM, covering diverse forms of implicit and explicit prompts (see Fig. 6 and Tab. 2). REALM3D can be used to evaluate the robustness of models across diverse applications. We provide details of REALM3D in the supplementary materials.
Baselines and metrics. We compare REALM against previous state-of-the-art methods for open-vocabulary 3D segmentation, including Gaga [21], GAGS [24], and GS-Group [35]. We report the mIoU and mBIoU following previous works  [31, 25, 13, 24] to quantitatively exanimate the accuracy of 3D reasoning segmentation results.
@@
Methods mIoU mBIoU
GS-Group (Baseline) 0.32 0.30
GS-Group+Qwen2.5-VL 0.78 0.77
+Global Reasoning 0.89 0.88
+Local Refinement 0.95 0.94
(a) Performance of each component in GLSpaG. Adding global reasoning and local refinement progressively improves performance over Qwen2.5-VL.
@@
(b) Global camera sampling. K-means view clustering and Top-K ID selection play a crucial role in the global camera sampling process.
@@
(d) K-means Clusters. Both insufficient and excessive clustering can harm multi-view reasoning.
Value of $N^{\text{global}}$ mIoU mBIoU
$N^{\text{global}}=4$ 0.81 0.80
$N^{\text{global}}=8$ 0.95 0.94
$N^{\text{global}}=16$ 0.95 0.94
(e) Number of global cameras. Our method is robust to the hyper-parameter $N^{\text{global}}$.
@@
(f) Local refinement steps. Excessive finetuning can lead to overfitting and degradation.
@@ 4.2 Main Results
Qualitative Comparisons. Previous methods enable 3D localization by leveraging the language understanding capabilities of CLIP [26] or Grounded-SAM [28]. While these approaches offer basic open-vocabulary querying capabilities, they lack the ability to perform reasoning over implicit instructions. We visualize the performance between REALM and baselines under different type of implicit queries. The results are presented in Fig. 5. Demos can be found in the supplementary materials.
@@
Quantitative Comparisons. We quantitatively evaluate the performance of REALM on both implicit and explicit queries. A subset of results is presented in Tab. 1. More results can be found in the supplementary materials.
@@
Language-Driven 3D Editing. With accurate 3D object localization, REALM enables precise and fine-grained scene editing without disturbing surrounding objects. As shown in Fig. 7, our model supports a variety of 3D editing tasks, including object removal, replacement, and stylization. REALM modifies the scene without interfering with adjacent content, ensuring faithful preservation of occlusion relationships. In tasks involving large-scale appearance changes, such as stylization, REALM effectively isolates the target object, leaving surrounding regions unaffected.
@@ 4.3 Ablation Study
Each component of GLSpaG. As shown in Tab. 3(a), we evaluate the performance after completing each stage of GLSpaG. Baseline method (GS-Group) is unable to handle implicit queries. When combined with Qwen2.5-VL, GS-Group can achieve substantial performance improvement. Since result is highly unstable, we report the average score in the table. With our Global Reasoning module, the agent can automatically select appropriate viewpoints for segmentation, leading to a stable mIoU of around 0.89. With the additional Local Refinement module, the predicted masks exhibit well-aligned boundaries (see Fig. 8).
@@
K-means clusters. As shown in Tab. 3(d), the number of clusters ($N^{\text{cluster}}$) significantly affects grounding accuracy. We set $N^{\text{cluster}}{=}24$ in this work, as both too few and too many clusters hinder the selection of optimal global views.
Number of global cameras. As shown int Tab. 3(e), we experiment with different values of $N^{\text{global}}$ and observe that the model remains relatively robust across this range. In practice, we set $N^{\text{global}}=8$.
Local refinement steps. Tab. 3(f) illustrates how the model’s performance evolves with an increasing number of local refinement steps. Since refinement is performed on a limited number of views, excessive optimization can lead to overfitting the 3D mask to those specific viewpoints, ultimately degrading the segmentation performance.